Please help
complete the ordered pairs so htat each is a soluction of the given linear equation. Then graph the equation.
equation x=5+y
X Y
0
-1
5

Answer by stanbon(75887)

complete the ordered pairs so htat each is a soluction of the given linear equation. Then graph the equation.
equation
x = y+5
y = x-5
--------------
X Y
0 ; y = 0-5 = -5
-------
-1; y = -1-5 = -6
-------
5 ; y = 5-5 = 0
===================
Cheers,
Stan H.